OpenStudy (anonymous):

Determine the zeros of f(x) = x^3 – 3x^2 – 16x + 48

OpenStudy (anonymous):

can you please go one step at a time

OpenStudy (anonymous):

Determine the zeros of f(x) = x^3 – 3x^2 – 16x + 48 The question asks you find the values of x for which f(x) is 0. Hence @callisto set \(f(x)=0 \)

OpenStudy (maheshmeghwal9):

yeah! \[f(x)=x^3-3x^2-16x+48=0.\]

OpenStudy (maheshmeghwal9):

The next step is taking common factors from first 2 & last 2.

OpenStudy (anonymous):

After that, the method employed is known as factorization of a cubic polynomial. It's basically just to factor out the common things and reduce it to a quadratic and then factor it to finally reduce it to three linear equation in product form. Finally, it's equating every term to zero and reducing the final answer :)

OpenStudy (maheshmeghwal9):

Yeah: ^_^

OpenStudy (anonymous):

To understand the last step Think about, \(a\times b = 0\), this can only happen when either of them is zero.

OpenStudy (anonymous):

The same is true about \( a\times b\times c =0 \)

OpenStudy (maheshmeghwal9):

@Callisto Why u deleted the solution? We were trying to make her understand ur solution:)

OpenStudy (maheshmeghwal9):

ur solution was best:)

OpenStudy (callisto):

It was not good at all....

OpenStudy (anonymous):

Hm but without your solution, our explanation are somewhat meaningless.

OpenStudy (maheshmeghwal9):

hmm........ i don't agree. becoz the second method is too hard for @sara1234 to understand.

OpenStudy (maheshmeghwal9):

so ur sol. was best @ all.

OpenStudy (anonymous):

okay so can we start by doing the rational root thrm

OpenStudy (anonymous):

Yes we can use rational root theorem. But it is a bit tedious compared to the solution posted by @callisto

OpenStudy (callisto):

Put f(x) =0 x^3 – 3x^2 – 16x + 48 = 0 x^2 (x-3) -16 ( x-3) = 0 (x^2-16) (x-3) =0 (x-4)(x+4)(x-3) =0 x+4 =0 or x-3 =0 or x-4=0 x = -4 or x =3 or x=4

OpenStudy (anonymous):

can we do it the other way I think it's easier for me to understand

OpenStudy (anonymous):

Factors of 48- +-1=+-2+-3+-4+-6+-8+-12+-16+-24+-48 Factors of x- +-1

OpenStudy (maheshmeghwal9):

We can find roots according to this too:) \[ax^3+bx^2+cx+d=x^3+(\alpha+\beta+\gamma)x+(\alpha \beta+\beta \gamma+\alpha \gamma)x+\alpha \beta \gamma.\]Where \[\color{blue}{\alpha, \beta, \gamma \space \text{are roots of this cubic eq..}}\]

OpenStudy (anonymous):

Using the rational root theorem, the possible roots are of the form: \[ \pm \frac{1, 2, 3, 4, 6, 8, 12, 16, 24, 48}{1} \] If you go through each of them you will see that that equation will yield zero at \(x=3,\pm4 \)

OpenStudy (anonymous):

okay so now we have to do rule of signs

OpenStudy (anonymous):

2 or 0 positive real zero

OpenStudy (anonymous):

right?

OpenStudy (maheshmeghwal9):

From my solution if we break \[\alpha \beta \gamma=+48.\]Such that \[\alpha+\beta+\gamma=-3.\]then we get \[\alpha=+4;\beta=-4;\gamma=-3\]So our roots will be \[- \alpha; -\beta ; -\gamma\]

OpenStudy (anonymous):

@maheshmeghwal9 i really don't understand what your trying to explain to me

OpenStudy (anonymous):

@maheshmeghwal9 's approach is also known as vieta's theorem.

OpenStudy (anonymous):

http://mathworld.wolfram.com/VietasTheorem.html

OpenStudy (maheshmeghwal9):

yeah of course:) I know that @im2bad no need to understand @sara1234 I was only showing another method

OpenStudy (maheshmeghwal9):

just concentrate on @Callisto 's method :D

OpenStudy (anonymous):

sorry im not that smart as you guys so can you dumb it up a bit lol

OpenStudy (callisto):

Have you learnt factorization ?

OpenStudy (anonymous):

the way i am trying to do this is rational thrm then rule of signs then fundamental thrm then quadratic formula but im having trouble

OpenStudy (callisto):

Hmmm you haven't answered my question :|

OpenStudy (theviper):

factorize & get ur zeroes

OpenStudy (ash2326):

@sara1234 did you understand ?

OpenStudy (anonymous):

Umm.... i'll just try another question

OpenStudy (ash2326):

No, I'll help you. Don't give up

OpenStudy (ash2326):

We'll use the rational factor theorem here, if there are any real roots, they will be of the forms \[\frac{\pm \ factors\ of\ 48}{1}=\frac{\pm 2. 2. 2.2.2 .3}{1}\] Do you get this?

OpenStudy (anonymous):

yes that's what i was trying to say we start of with

OpenStudy (ash2326):

Yeah, try +2 or -2, see if any of these two is a zero

OpenStudy (anonymous):

2 or 0 positive real zero

OpenStudy (ash2326):

That's correct, I meant that check if +2 or -2 is a zero of f(x) Check if f(2)=0 or f(-2)=0

OpenStudy (anonymous):

can we continue this tommorw i have to go i really apperciate your help

OpenStudy (ash2326):

Sure thing, tag me to call me :)